Suppose, 10 boxes of facemask are being sold on $50 each and due to covid-19, supply of facemasks increase. Due to the pandemic, price increased upto $60 and there is increase of quantity of 20 percent. Find the price elasticity of supply as well as explain which kind of elasticity is it?


1
Expert's answer
2021-05-18T11:09:04-0400

The price elasticity of supply is:

"Es = \frac{12 - 10} {60 - 50} ×\frac{60 + 50} {12 + 10} = 2/10×110/22 = 1."

So, the supply is unit-elastic.